mathslover (mathslover):

Evaluate \(\lim_{x\to 0} \sin^{-1} \left( \cfrac{\cos^{-1} x + \cos^{-1} (x^2) }{\pi} \right) \)

ganeshie8 (ganeshie8):

Hint : \[\lim f(g(x)) = f(\lim g(x))\] if \(f(x)\) is a continuous function

mathslover (mathslover):

I don't know, I'm getting the answer as 0 .. but the book suggests that it doesn't exist :/

OpenStudy (loser66):

Is it not that when we replace x = 0, we have cos^(-1) (0) = pi/2, hence lim sin^(-1) (1) = pi/2?

ganeshie8 (ganeshie8):

^

mathslover (mathslover):

Oh yeah, sorry for that. But, still, how does the limit not exist?

mathslover (mathslover):

And here is the reason it has provided: "Limit does not exist as LHL does not exist." \(\lim_{x \to 0^{-}} \left( \cfrac{\cos^{-1} x + \cos^{-1} (x^2)}{\pi} \right) = 1^{+} \)

ganeshie8 (ganeshie8):

book is wrong

mathslover (mathslover):

Oh okay. Thanks a lot bhaiya! :)

ganeshie8 (ganeshie8):

wait a second, maybe book is right... what is the domain of arccos(x) ?

mathslover (mathslover):

It is -1 to 1 [-1,1]

ganeshie8 (ganeshie8):

Ahh right, we're good then. Limit exists and equals pi./2 :)

mathslover (mathslover):

great, thanks again bhaiya

ganeshie8 (ganeshie8):

\[\large \lim\limits_{x\to 1} ~\arcsin(x) \stackrel{?}{ =}\frac{\pi}{2}\]

mathslover (mathslover):

Yeah! I got it bhaiya...

ganeshie8 (ganeshie8):

that is a trick question actually, notice that x=1 is a boundary point arcsinx is defined in [-1, 1]

mathslover (mathslover):

Oh, but that will disturb the differentiability of the function, right? How's it going to affect the limit?

ganeshie8 (ganeshie8):

You're probably right, can you give me a convincing argument why the limit exists (recall that both side limits must exist and be same for the overall limit to exist)

mathslover (mathslover):

I see what you mean to point here. I'm just confused that we have arcsin(1) here, right? Then why will we take case of \(1^{+}\) here?

mathslover (mathslover):

Even if we are considering RHL, then that doesn't mean we can simply put 1(+) at the place of 1 there...! right? And the book shows that LHL is arcsin(1(+))

ganeshie8 (ganeshie8):

we simply don't worry about RHL here because the function itself is not defined for \(x\gt 1\) a sequence can converge happily to a boundary point, no issue with it i guess @zzr0ck3r @eliassaab

OpenStudy (anonymous):

then why don't we do the same for \[\sqrt{x} \] ?

mathslover (mathslover):

Yeah, but here, x is tending to zero. Why will we take case of x >1 ? :/ Sorry if these questions sound stupid but..

ganeshie8 (ganeshie8):

@Catch.me are you suggesting \(\lim\limits_{x\to 0^{}}~\sqrt{x}\) doesn't exist if we think of \(\sqrt{x}\) as a real valued function ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so I can't understand your reasoning!

ganeshie8 (ganeshie8):

Okay :) may i knw why do you think \(\lim\limits_{x\to 0^{}}~\sqrt{x}\) doesn't exist ?

mathslover (mathslover):

May be because LHL doesn't exist?

OpenStudy (misty1212):

limit exists if \[\lim_{x\to a^+}f(x)=\lim_{x\to a^-}f(x)=L\]